The Traffic Police department is concerned with the increase in Accidents on Shaheed-Millat Expressway.They are setting up a dedicated team to analyse the causes and propose a solution. Accidents can occur because of the sharp turns on road or unavailability of Pedistrian bridges. On the other hand, accidents can occur due to unsafe driving or human error. The observation was further breakdowned in


Shift Unsafe Road Conditions Human Error Morning 6% 28% Day 4% 12% Evening 5% 26% Night 3% 16% Table 5

If an accident report is selected randomly from the 170 reports, (a) what is the probability that the accident occurred on the Evening shift? (b) what is the probability that the accident occurred due to human error? (c) what is the probability that the accident occurred due to unsafe road conditions? (d) what is the probability that the accident occurred on either the evening or the night shift?
1
Expert's answer
2020-06-15T18:28:31-0400

a) P=(5+26)/100=0.31

b)P=(28+12+26+16)/100=0.82

c)P=(6+4+5+3)/100=0.18

d)P=(5+26)/100+(3+16)/100=0.5